A constant volume gas thermometer records pressure of body as 250mm of Hg at 0°c and 350mm Hg at 100°c. Calculate the temperature when the gas pressure reads 300mm Hg
1
Expert's answer
2020-03-10T11:36:42-0400
"T=aP+b"

"0=250a+b"

"100=350a+b"

So,


"100=100a\to a=1"

"0=250+b\to b=-250"

Then,


"T=P-250"

"T(300)=300-250=50\degree C"